A square board that has an area of 25 square inches is to be cut into pieces, each of which is a square with sides of length 1,2 or 3 inches. What is the least number of such square pieces into which the board can be cut?
Answer by edjones(7311)

Obviously the more 3" blocks the better, MAYBE!
Try 8x8 3" square blocks=24^2=576"^2 Then we need 625-576=49 1" sq blocks for a total of 64+49=113 blocks.
Try 7x7 3" sq blocks=21^2=441"^2 Then we need 625-441=184"^2 of 2" sq blocks
184/4=46 blocks. 49+46=95 blocks. Answer.
.
Ed